OpenStudy (anonymous):

Is the equation true, false, or open? Explain. 9p + 8 = 10p + 7

OpenStudy (anonymous):

if p=1 then the statement is true

OpenStudy (stormfire1):

9p+8=10p+7 9p-10p = 7-8 -1p = -1 p=1...so if p =1 it's true